OpenStudy (anonymous):

anyone know about summation notations? How do you get the number your dividing by?

OpenStudy (amistre64):

a little more information would be helpful

OpenStudy (amistre64):

if this is part of a problem, knowing the problem would be useful :)

OpenStudy (across):

\[\sum\]This is sigma, a pretty cute symbol standing for summation. For some reason, he's obsessed with infinity, though.

OpenStudy (anonymous):

what do you need? \[(X-X)^2 for the following values of X: 5,7, 2, 1, 3, 6\]

OpenStudy (amistre64):

5-5=0 7-7=0 .....

OpenStudy (anonymous):

what do you put on the bottom of the equation? How do you know what number to use

OpenStudy (across):

As Jack Sparrow would say: "You're not making any sense, mate."

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

if your talking about the index, it tends to be the 1

OpenStudy (amistre64):

\[\sum_{i=1}^{n}(x_i-x)^2\]maybe?

OpenStudy (amistre64):

the top number is the number of elements you have in the sequence

OpenStudy (amistre64):

the bottom tends to tell you where to start at

OpenStudy (amistre64):

and sum(x^2) = n(n+1)(n+(n+1))/6

OpenStudy (across):

\[\sum_{i=1}^{n}x_{i}=x_{1}+x_{2}+...+x_{n}=\frac{1}{2}n(n+1)\]

OpenStudy (amistre64):

x^3 is the square of x ;)

OpenStudy (anonymous):

Sigma(X-X)^2 for the following values of 5,7,2,1,3,6 Note: The second X has an -over the top of it. Can anyone tell me how you know what you use on the bottom of the division

OpenStudy (amistre64):

x bar tends to be notation for mean

OpenStudy (amistre64):

add up your data set and divide by ... there is 6 of them so .. divide it by 6 to get the valueof x bar

OpenStudy (amistre64):

if your going for variance; it depends on sample or population

OpenStudy (amistre64):

the data set here suggests a population since it doesnt state otherwise

OpenStudy (anonymous):

Oh okay. That's exactly wahat I was needing to know. Thank you so so much!

OpenStudy (amistre64):

for a sample, we use a degree of freedom, usually n-1

OpenStudy (amistre64):

for the population; just n :) youre welcome

OpenStudy (amistre64):

also of note: \(\bar x\) is usually used for sample mean whereas, \(\mu\) is used to indicate a population mean

OpenStudy (amistre64):

sample tends to look like this \[\frac{\sum(x-\bar x)^2}{n-1}\] and population like this \[\frac{\sum(x-\mu x)^2}{N}\]

OpenStudy (amistre64):

typoed that last bit .... but your smart enough to read thru it